One pump can fill a water tank in 40 minutes and another pump takes 30 minutes. When the tank was empty, both pumps were turned on for 10 minutes and then the slower pump was turned off. How much longer did the faster pump have to run before the tank was filled?
Answer by stanbon(75887)

One pump can fill a water tank in 40 minutes and another pump takes 30 minutes. When the tank was empty, both pumps were turned on for 10 minutes and then the slower pump was turned off. How much longer did the faster pump have to run before the tank was filled?
--------------------------------
1st pump rate = 1/40 job/min
---
2nd pump rate = 1/30 job/min
---
Together rate = 1/40 + 1/30 = (70)/1200 = 7/120 job/min
==============================
Equation:
work done + work done = one job done
(7/120)10 + (1/30)x = 1 job
Multiply thru by 120 to get:
70 + 4x = 120
4x = 50
x = 12.5 minutes (time the faster pump had to run)
============
Cheers,
Stan H.